PHOTOCHEMISTRY OF MATRIX-ISOLATED VINYL ACETATE

KRISTA COHEN and \underline{C. A. BAUMANN}, Department of Chemistry, The University of Scranton, Scranton, PA 18510-4626.

Vinyl acetate (C_{4}H_{6}O_{2}) was isolated in matrices of nitrogen, argon, and krypton. The ultraviolet irradiation (λ< 250 nm) of the matrix results in the formation of a number of products including carbon dioxide, carbon monoxide and ketene (CH_{2}CO). Irradiation of the vinyl-^{13}C_{2} labeled compound yields a mixture of labeled and unlabeled CH_{2}CO, CO_{2} and CO, indicating that these products arise from the acetyl and the vinoxy moieties. The product ^{12}C/^{13}C ratios vary with matrix (N_{2}>Ar>Kr). The photolysis efficiency is reduced when the parent compound is deuterated.